OpenStudy (anonymous):

Find a rational number and an irrational number that are between 5.2 and 5.5. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (anonymous):

this si the last question.

OpenStudy (michele_laino):

we can write this: \[\Large 5.2 = \frac{{52}}{{10}},\quad 5.5 = \frac{{55}}{{10}}\]

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

so a rational number between 5.2 and 5.5 can be: \[\Large \frac{{54}}{{10}} = 5.4\]

OpenStudy (michele_laino):

since we have: \[\Large \frac{{52}}{{10}} < \frac{{54}}{{10}} < \frac{{55}}{{10}}\]

OpenStudy (anonymous):

so 5.4 is the answer

OpenStudy (michele_laino):

yes it is the first part of your answer, now we have to find an irrational number

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

an irrational number is pi=3.14159...

OpenStudy (michele_laino):

nevertheless 3.14159...<5.2

OpenStudy (michele_laino):

so we consider this number: 1.7* pi, namely: 1.7*3.14159... now, what is 1.7*3.14159..=...?

OpenStudy (anonymous):

um

OpenStudy (michele_laino):

namely, what is: \[\Large 1.7 \cdot \pi = ...?\]

OpenStudy (michele_laino):

you can use your calculator

OpenStudy (michele_laino):

I got this: 5.3407075111026485053864937515752... am I right?

OpenStudy (anonymous):

yes thats what i got

OpenStudy (michele_laino):

and that number is greater than 5.2 and less than 5.5, so it is the requested irrational number

OpenStudy (michele_laino):

we can rewrite it like below: \[\Large 1.7 \cdot \pi = \frac{{17\pi }}{{10}}\]

OpenStudy (michele_laino):

and, as you can see, its approximated value to the nearest hundredth is 5.34

OpenStudy (michele_laino):

and we can write this: \[\Large 5.2 < \frac{{17\pi }}{{10}} < 5.5\]

OpenStudy (anonymous):

sorry i had to go eat

OpenStudy (anonymous):

um so 17/10 is the answer for irrational

OpenStudy (michele_laino):

it is: \[\Large \frac{{17\pi }}{{10}}\] please study my replies here

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

:)